OpenStudy (markayla01):

Write the trigonometric expression as an algebraic expression. 2 cos(2 arccos x)

OpenStudy (anonymous):

use \(\cos(2\theta) =\cos^2(\theta) - \sin^2(\theta) = 2\cos^2(\theta) - 1\)

OpenStudy (anonymous):

Then \[ 2\cos(2\arccos(x)) = 2(2\cos^2(\arccos(x)) - 1) \]

OpenStudy (anonymous):

In this case \(\theta = \arccos(x)\)